MOBY-L267 experiment with MOBYrefresh - BS04cfg001-004 MOBY264 precals testing

Mike checked the length of the screws on BS04 and they looked OK.

BS04cfg001 = pre-L264, 29No-4 Dec  = bare fibers
BS04cfg002 = pre-L264, Dec 14      = shutters installed (see Log Sheet for serial numbers)
BS04cfg003 = pre-L264, Dec 15      = Mark Iris removed from Track 1 and Mike tightened down the locking screws for the 4x irises I adjusted   yesterday
BS04cfg004 = pre-L264, Dec 16      = added the "window plate" to the BS shutter block, then coupled the window plate to the FO block in the tent
Problems seen (and solved) and good things:
  1. Mirrored UV ghosting of light from illuminated tracks (Page Num 9.04). Also seen on in-water data.

Log Sheets

Photos

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 29 Nov Raw Data , Table
DAY01 - Stability testing (first day) - all track OL455
29 Nov 2017
1.02
Day 02 - 30 Nov Raw Data , Table
DAY02 - Stability testing (second day - no movement) - all track OL455
29 Nov 2017
1.03
Day 03 - 1 Dec Raw Data , Table
DAY03 - Stability testing (third day - no movement) - all track OL455
30 Nov 2017
1.04
Day 04 - 1 Dec Raw Data , Table
DAY04 - Stability testing (fourth day - no movement) - all track OL455
1 Dec 2017
1.05
Day 05 - 4 Dec Raw Data , Table
DAY05 - Stability testing (fifth day - no movement) - all track OL455
4 Dec 2017
CFG002: 14 iris and 14x shutters installed, The 4x Es/Ed tracks also have a 1.0 ND filter installed. he BS04cfg002 trks are:  2=Es;  5,6=Ed,LuT;  8,9=Ed,LuM; 12,13=Ed,LuB, but we will not have Ed collectors on the M264 buoy.
1.06
Day 06 - 14 Dec Raw Data , Table
DAY06 - Iris and neutral density filter track reduction check
14 Dec 2017
CFG003: Mark Iris removed from Track 1 and Mike tightened down the locking screws for the 4x irises I adjusted yesterday
1.07
Day 07 - 15 Dec Raw Data , Table
DAY07 - recheck iris reductions and it time cal
14 Dec 2017
CFG004: added the "window plate" to the BS shutter block, then coupled the window plate to the FO block in the tent
1.08
Day 08 - 16 Dec Raw Data , Table
DAY08 - system responses
14 Dec 2017
 
Day01 - stability
2.01 find tracks Find tracks from day 1 (redefining the tracks to the full track data) 29 Nov 2017
Day06 - Iris and neutral density filter track reduction check
7.01 Iris reduction Iris check for Es, Ed and Lu (iris -> iris adj) 14 Dec 2017
7.02 Iris reduction V2 Iris check for Es, Ed and Lu (rel to day 5 - bare fiber -> iris adj) 14 Dec 2017
Day07 - recheck iris reductions and it time cal
8.01 saddle shapes Saddle shape changes from day 5- 7 15 Dec 2017
8.02 int time Int time cal (not including dimmer) 15 Dec 2017
8.03 int time w/ dimmer Int time cal (including dimmer) 21 May 2019
8.04 int time w/ dimmer - 10 sec Int time cal (including dimmer) - 10 sec normalization 21 May 2019
8.05 compare 1 and 10 sec cal Comparing the 1 and 10 sec normalize int time cal 21 May 2019
Day08 - System response
9.01 saddle shapes Saddle shape changes from day 5- 8 (day 8 knife shapes appear) 15 Dec 2017
9.02 Lu system rsps Lu system responses 18 Dec 2017
9.03 Es system rsps Es system responses 18 Dec 2017
9.04 UV ghosting Ghosting of UV data from other tracks 18 Jan 2018
9.05 cross trk ratio Cross track ratios 13 Mar 2018
Compilation of day data
20.01 UV nets->BS link RS pow lix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here 31 Jan 2019
Multi Day graphs (cfg001 data - bar fibers - no shutters installed)
100.01 Comparing track movement Comparing track movement for stability data  
100.02 Stability Looking at the lamp stability @ 3 pixels  
100.03 Spectral Stability Looking at the lamp with ratios @ 3 pixels  
100.04 Background Stability Looking at the backgrounds stability @ 3 pixels  
100.05 Dark Stability Looking at the dark stability @ 3 pixels  
100.06 Dark and averaging Looking at darks and averaging the data  
100.07 Background-Darks Mean of 24 Backgrounds - Mean of 10 Darks  
100.08 Aux data Looking at Auxilliary data  
100.09 Hot Pixel Hot pixel stability  
100.10 Track to Track stability Track to track stability (ratio each track to track 1)  
100.11 Track movement Track movment effect with all tracks on and full track definitions  
100.12 UV data stability Looking at the nets in the UV region  
100.13 looking at bac/lite var on UV area Looking at the bac variability relative to lite in the UV  
100.14 Ratio day6/day7 Looking at the bac variability relative to lite in the UV  

 

DAY 1 EMAIL from Mike

Hello Stephanie,
Are you ready for some more BS data?
Today was Day01 of pre-MOBY264 BS04cfg001 stability testing.
I am calling this cruise L267 (because the last MOBY263 dive cal was L266).

There are photos of the setup at /ftp1/Mike/L267/bs4cal/photos/,
which is the the BS04 bare FO-bundle fibers looking into the NIST OL455 sphere.

The first few photos show FO#13 which had some epoxy gooped around the
screw-down sleeve of the FC connector, which I tried to pick/cut out so the
FC connector could be connected.

The idea was to run this setup for a week to see if it remains stable.
My method today was to scan 4x backgrounds, then 5x signals, then 4x bac's,
and repeat 5x times - kinda like a MOBY profile.  At the end, after the lamp
was turned off I scanned 10x darks.  The exposure time was 5 sec for all scans.
The total was 25x sig scans + 24 bac scans + 10 drk scans.
The .fits files are at /ftp1/Mike/L267/bs4cal/day01/.

If you have a better idea for this setup, we can change it for tomorrow...

I recorded the OL455 photodiode detector output (Amps) at
/ftp1/Mike/L267/bs4cal/day01/2017112901-ol455log.txt
I am sorry to say that I made one change to this file format,
by adding the date to the time stamp!

The OL455 seemed stable-ish over the 1/2 hour of this run -
aboot 0.3% PD Amp change vs mean - which seemed OK-ish at the
very low light level needed to get a reasonably-long exposure time.

I recorded the Aux data as /ftp1/Mike/L267/bs4cal/day01/2017112901-LJlog.dat
which should be the internal Resonon thermistors & relative humidity,
plus external/ambient %RH + degC, and the extra degC thermistor in air
(the last degC is available to screw down to the splitter, for example) -
there are 2 photos of the aux data cabling & LabJack setup.

There is one log sheet at /ftp1/Mike/L267/bs4cal/doc/.
I think that is aboot it.  Easy right?
Thank you and Good Luck!  MF

DAY 2 EMAIL from Mike

HI StePh,
Day02 data are on the NAS =
same routine as day01, without any new photos,
and, it seems that the operator failed to save file s...13.fits - sowee...
MF

EMAIL from Mike on the BS04 screws

Hi Steph,
FYI: Today I opened the BS04 to check the length of internal screws.
All securing screws were glued down, so I did not remove any of them.
There are some new photos of this at /ftp1/Mike/L267/bs4cal/photos.

And, Mark gave I a list of shutter serial numbers for the BS04cfg001+,
/ftp1/Mike/L267/bs4cal/doc/20171203_BS04cfg002_shutter_SN's.jpg

MF

EMAIL from Mike on the BS04 screws

Mark thought I should check the length of the screws inside the BS4,
so I removed two, and they were OK - there are 2x more photos of this on the NAS...

Day 6 EMAIL from Mike

Hi Steph,

The BS04 got a shutter block and shutter controller installed today,
so it is now called BS04cfg002.

There is an iris with each of the 14x shutters.
I think we maybe should have put irises only at 7x ports (Es+3xEd+3xLu) (?),
because there was no "non-attenuated" track to use as a reference whilst
trying to adjust the irises.

The 4x Es/Ed tracks also have a 1.0 ND filter installed.

I tried to adjust the 3x Lu tracks net-signal down 10x via the irises.
Note that this will be a greater-than-10x-decrease versus a "bare" FO,
because the 10x decrease is relative to the open iris net-signal...

I tried to adjust the Es track net-signal down 0.7x via the iris.
I came up with this via your plot s2017101506_img.mat,
where it looked like the M263/BS03cfg03 Es signal(ADU) was down ~2.5x
versus the other Ed & Lu signals, and since we did not include the extra
0.5ND filter on M264/BS04cfg002, which should give ~3.2x > signal,
3.2 - 2.5 = 0.7x.

The BS04cfg002 trks are:  2=Es;  5,6=Ed,LuT;  8,9=Ed,LuM; 12,13=Ed,LuB,
but we will not have Ed collectors on the M264 buoy.

The day06 *.fits files are at /ftp1/Mike/L267/bs4cal/day06/
There are photos at /ftp1/Mike/L267/bs4cal/photos/20171213*.jpg
And 1x log sheet + 2x green-pad-note scans at .../doc/*day06.jpg

Aloha, MF

Day 7 EMAIL from Mike

Hi Steph,
I do here want first to apologize for the mess I made below.

Today, after the telecon, Mark removed the iris from trk#1,
and I tightened down the locking screws for the 4x irises I adjusted yesterday,
so I am calling this BS04cfg003.

I had to move the BS back from the OL455 sphere to get at the lock screws,
so the setup changed from yesterday day06, but otherwise it is the same setup.

Tonight / day07 I logged the OL455 PD monitor to 2017121501-ol455log.txt
and I acquired Aux data to 2017121501-LJlog.dat.  There are 2x log sheets.

I started out scanning 5x bac,sig pairs with all shutters open, to look at the
difference between Trk#1 without an iris and the other Trk's with iris,
and compare this with day06 scans with iris @ T1.
Then, I closed the BS shutters and collected 5x dark scans (lamp = on).
These are files named %2017121501:05.fits

Then I tried to get an exposure time calibration data set via Trk#1 (only T1 shtr open).
This was supposed to follow the same acq approach of post-M262/BS01, 01-Oct-2017, post-L264,
with the addition of some longer exp times at the end to go from 100 to 300 sec.

The exp file names have a 3-digit sequence number: %20171215001...036.fits
- not 2-digit seq num like the above files

However, I ran into some trouble with the camera locking up.
It "stalled" 3x times, and maybe at the 4th time I setup a fan blowing air
into the cooling intake on the side of the camera.  After the fan was on
the camera did not hickup again, so hopefully it was just a camera
overheating problem...  So, it would be nice to look carefully at the
Aux data to see if there is any sign of overheating at the internal thermistors.
The place/times of the camera stalls are marked with red stars on the
log sheets.  There are 3x red stars, but I think there was a 4th lockup...

On pg.2 of the log sheet, at the 3rd or 4th camera stall I got the file numbering
out-of-order.  I messed up somewhere before seq # 024.  I think when "we"
look at the exp times it will be obvious where I messed up.
I think seq# 024:027 are OK.  I re-ran the questionable prior-set as seq# 028:036.
This will make patching this run together a challenge...

I'm thinking a first-order check might be just to normalize everything to
the PD monitor Amps, and see if there are any obvious flyers.
I'm concerned because it would be nice to determine if the camera-stall
problems caused any data problems.  This would potentially impact whether
or not this camera is deployable in MOBY264 (yikes!)...

If this exposure cal seems to make sense then we will probably move
the BS out to the tent tomorrow and continue with Es and Lu cals...
Thank You, MF

Day 8 EMAIL from Mike

Hi Steph,
Today I added the "window plate" to the BS shutter block,
then coupled the window plate to the FO block in the tent,
so this would be cfg004 for the BS04.

Then the Es sensor scanned the FEL-F600(C3),
and the 3x Lu's scanned the OL525(L10),
which would be day08 data, which had 2x log sheets.
I took some photos today, but I have not FTPed them yet.

The OL425's controller - the OL450 - had problems during the
last run for the LuTop, but I think the BS data should be OK.
It looks like I failed to collect the 3x darks after the LuMid lites -
the files d2017121616:18.fits are missing - sorry again.

I don't know if you have the new OL425(L10) cal yet -
if not, I am attaching it here.
Aloha, MF

<attached: OL425L10.txt>

PS email:

I meant to say:
I did not get to scan wave cal lamps with the BS04,
so we'll have to rely on your Resonon cal plus
Fraunhoffer lines ... from (hopefully) next week.  The plan
is to finish assembling the BS tomorrow and get it
into the bottom instrument bay on MOBY.
If there is a chance I will try to get wave lamp scans
and/or shine some sunlight on Es...